OpenStudy (anonymous):

PLEASE HELP me X-box factor. please asap. 29.) -7x^2-54x+16=0 30.) 24x^2-52x+20=0 32.) 15x^2-2x-1=0 33.) 14x^2+33x-5=0 34.) -15x^2+7x+4=0 PLEASE HELP me x-box factor a few of these??????

OpenStudy (primeralph):

X-Box?

OpenStudy (anonymous):

yeah X-box factor where u have 4 boxes and a X. I need help solving these problems like that @primeralph . Please help?

OpenStudy (primeralph):

I don't know that. The X-Box I know is very different.

OpenStudy (campbell_st):

well its not x box... but its a great method that works for \[ax^2 + bx + c \] multiply a and c in your question 1 its -7 x 16 = - 112 find the factors of -112 that add to b -54 the larger is negative so its -56 and 2 then write the binomials as (ax + factor 1)(ax + factor 2) -------------------------- a factor the binomials and your find a common(2) factor which the denominator that cancel so in your question its (-7x - 56)(-7x + 2) -7(x + 8)(-7x + 2) ------------------ = ------------------ = 0 -7 -7 cancel the common factors and you are left with (x + 8)(-7x + 2) = 0 now you can solve the binomials for the solution

OpenStudy (anonymous):

thanks

OpenStudy (campbell_st):

question 30 divide every term by 4 1st and you get \[6x^2 - 13x + 5 = 0\] then follow the steps above.

OpenStudy (campbell_st):

in 30 both factors are negative...